# Panoptic: True Joint mmWave Communication and Sensing with Compressive Sidelobe Forming

### v1

@@ IV Implementation
Hardware. We use commercial mmWave phased arrays as front-ends at both TX and RX. Specifically, we use Sivers EVK06002 boards [40] that can operate in 57-71 GHz regime and are widely employed in previous mmWave research and development studies. These boards have antennas for transmission and reception separately, as shown in Fig.4(a). The array supports beam scanning in the azimuthal direction with the field of view of 90, i.e., the mainlobe spans from -45 to 45 with a step size of about 1.5. We generate a compressive sidelobe beambook for any given mainlobe direction by randomly turning off a subset of antennas without modifying the phase shifters’ values. If not specified, we turn off four out of sixteen antennas, and we consider different random subsets, i.e., sidelobe configurations. We evaluate the impact of antenna subset size and number of observations in Sec. V-C.
@@
We have implemented Panoptic using both benchtop equipment as well as high-speed RFSoCs. In the former, we use an arbitrary waveform generator and real-time scope as baseband units on the TX and RX sides, respectively. In this case, the symbol rate is MHz and our frame structure contains 20 preamble symbols followed by 1500 data symbols. We also implemented Panoptic programmable RFSoC FPGA as baseband units on both ends. The ADC sample rate is GHz, and the bandwidth is GHz with 192 active subcarriers. Our frame structure contains a standard IEEE 802.11 preamble and 64 data symbols.
On the TX side, the mmWave array is connected to the baseband unit, sending modulated data streams with different modulation and coding rates. On the RX side, the downconverted IQ samples are captured and processed by another baseband unit. Differential IQ ports are connected for both TX and RX to remove the distortion and noise. The TX and RX clocks are not synchronized (the nodes are not wire-connected). Panoptic’s algorithm is written in a single Python code controlling the Sivers array configuration and data capture. For all experiments, the center frequency is GHz. If not specified, we place the TX and RX 3 m apart (we evaluate the impact of TX-RX distance in Sec. V-D). The preambles are used for coarse carrier frequency offset removal adopting conventional signal processing schemes from the literature [41]. The data symbols are used for joint data recovery and NLOS sensing.
@@ V-A Compressive Sidelobe Radiation Patterns
Results. We measure the radiation pattern under different mainlobe directions and we randomly turn off 4 out of 16 antenna elements. The measured radiation patterns are shown in Fig. 5. We show examples of two different mainlobe configurations and two distinct random sidelobe patterns under each mainlobe. In addition, we have simulated the theoretical radiation pattern using the knowledge of antenna geometry, inter-element spacing, and randomly chosen antenna subsets in measurements. We note that the internal nonidealities between antenna elements (e.g., due to variations in the lengths of the transmission lines in the circuit board connecting the RFIC to the antennas), require a one-time calibration for the array. We have performed antenna calibration based on the beambook provided by the manufacturer [42]. Fig. 5 illustrates good agreement between the simulated and measured radiation patterns. Note that we collected tens of different radiation patterns overall and validated the random sidelobe generation in all cases. Fig. 5(c) and Fig. 5(d) depict a few examples of measured sidelobe-varying beam patterns while the mainlobe remains unchanged toward 0 and 45, respectively.
@@ V-C2 Number of data symbols per beam ()
Second, Fig.7(b) shows the impact of the number of data symbols captured per compressive beam at the receiver. Recall from Eq. (8) that several measurements with a fixed beam help with averaging out noise and better mapping the sidelobe perturbations to the NLOS path signature. This is also reflected in our experimental results in Fig.7(b). We observe that the mean sensing error decreases with more observations per beam. With 25 symbols per beam, the mean NLOS angle error is 2.57, very close to the final value. The average reflector estimation error decreases below 2 after the symbol number exceeds 700. Also, the standard deviation decreases with more observation which indicates a robust estimation.
@@ V-C3 Antenna subset Size ()
Finally, we evaluate the impact of antenna subset size, i.e., the number of randomly selected OFF antennas. To this end, we repeat the experiments by turning off 1,2,3, and 4 antennas (out of 16) while placing a reflector at various angular locations within the range of -30 to 30 degrees (relative to TX). Recall that when all antennas are ON, no sidelobe perturbation happens, and hence NLOS sensing is not possible. We note that the total number of possible sidelobe configurations is only = =16 and = =120, for the case with 1 and 2 antennas OFF, respectively. For a larger number of OFF antennas, the random beam space size is much larger than the beambook memory in the Sivers arrays and hence we limit the number of tested beam configurations to .
@@
On the other side, when turning off more antennas, the mainlobe directivity gain is reduced, which is also captured in Fig.8(b) (all antennas ON in the leftmost). This may hinder the data detection performance with Panoptic. In this particular experiment, the BER remained at 0 (out of transmitted bits) in all cases as the gain of the mainlobe was sufficiently high to close a 3-meter link between the TX and RX even when using 12 antennas only. We evaluate the impact of TX-RX distance in Sec. V-D. We underline that the communication penalty becomes negligible with a larger antenna array because fewer antennas have to be turned OFF. For example, the randomness space size for the 16-element array with 4 antennas OFF is , and the size for a 64-element array with 2 antennas OFF is . However, turning off 2 out of 64 antennas decreases the mainlobe directivity (and consequently the SNR of the data link) roughly by dB. According to the MCS -EVM table in IEEE 802.11ad MCS [23], such an SNR reduction will not trigger an MCS downgrade. Therefore, Panoptic can offer the same achievable data rate and EVM performance while enabling seamless and real-time reflector sensing capability. In practice, the antenna subset size can be adapted in Panoptic to balance between angular localization accuracy and the desired BER.
@@ V-D4 Mobile Reflector Trajectory
So far, we have configured a reflector moving perpendicular to the TX-RX LOS path, similar to Trace 1 in Fig. 12(a). Here, we argue that Panoptic is robust to other arbitrary trajectories that a reflector might take as long as the object’s reflection footprint can be observed at the receiver. Hence, we place the TX and RX at a 3 m distance in a conference room and transmit 4QAM data symbols at a rate of 25 MHz. We configure three different trajectories as shown in Fig. 12(a). The estimation results are depicted in Fig. 12(b), 12(c) and 12(d), respectively. We observe that the ground truth angles match well with angles estimated by Panoptic. The error is higher around 0 for all three traces as the reflector is partially blocking the LOS path. Further, the error increases as the reflector approaches the FoV of TX or RX antenna arrays. Overall, we observe that the NLOS sensing accuracy is oblivious to the distance of the reflector from the TX or RX.
@@
Finally, it is worth mentioning that while we adopt a relatively low antenna switching rate for proof of concept experiments, it is possible to increase the antenna switching rate in practice for faster environmental tracking. With a simple calculation, we find that the antenna switching rate needed to angularly track mobile targets is well below the capability of today’s off-the-shelf arrays, allowing for real-time estimation and angular tracking in practical mobile environments. Specifically, given a symbol rate of and the number of data symbols per beam , the beam switching rate can be found as . For instance, with a symbol rate of GHz and minimum required (for reliable NLOS signature extraction as discussed in Sec. V-D), we reach the beam switching rate of MHz, which is an order of magnitude smaller than the Sivers array switching rate of MHz. Indeed, previous work showed an antenna switching rate of about MHz on Sivers arrays with FPGA [15]. Note that supporting higher symbol rates is trivial as one can consider increasing , which helps with better averaging and NLOS signature extraction from weak sidelobe perturbations. Further, as shown by our experimental results in Sec. V-D, testing about different random sidelobe configurations is sufficient for accurate NLOS sensing. Hence, the sensing frequency is calculated by KHz, following the previous example. Such a high sensing rate is sufficient to angularly track even high-speed targets, e.g., cars moving at the velocity of 100 km/h.
@@ VI Discussion
Super-Resolution Sensing with Panoptic. While this paper builds a novel foundation for joint communication and sensing using antenna subset switching, there are several schemes that can be adopted in future work to improve the sensing resolution. First, combining phase and amplitude readings, one can extend Panoptic to more advanced sensing tasks like environmental imaging, human activity recognition, etc. We leave this exploration for future work. Second, Panoptic chooses the antenna subset randomly without any feedback from the RX. In principle, not every sidelobe perturbation pattern provides the same amount of information for a particular target extraction. In the future, we will explore adaptive compressive sidelobe selection strategies. Last but not least, more advanced data processing methods such as data-driven and ML-assisted techniques could potentially improve sensing accuracy. We leave this for future work.
Scaling to Large Arrays . The mmWave nodes we use in evaluating Panoptic have 16 antenna elements on the azimuth plane. Indeed, mmWave AP/BS (and even mmWave handheld devices) are often equipped with larger antenna arrays. Larger antenna arrays expand the randomness space for sidelobe perturbations and result in better angular localization resolutions. In this case, turning off a small number of antennas can provide sufficient random space for sensing; hence, the penalty of compressive sidelobe transmission on the communication SNR and BER becomes negligible. As discussed in Sec. III-C2, the number of compressive sidelobe beam measurements only grows with ; hence, Panoptic is scalable to large arrays.

### v2

@@ IV Implementation
Hardware. We use commercial mmWave phased arrays as front-ends at both TX and RX. Specifically, we use Sivers EVK06002 boards [42] that can operate in 57-71 GHz regime and are widely employed in previous mmWave research and development studies. These boards have antennas for transmission and reception separately, as shown in Fig.4(a). The array supports beam scanning in the azimuthal direction with the field of view of 90, i.e., the mainlobe spans from -45 to 45 with a step size of about 1.5. The Sivers hardware has a maximum beam switching rate over 50 MHz from its datasheet. We generate a compressive sidelobe beambook for any given mainlobe direction by randomly turning off a subset of antennas without modifying the phase shifters’ values. We randomly chose a random subset of antennas to configure each, without any selection optimization. If not specified, we turn off four out of sixteen antennas, and we consider different random subsets, i.e., sidelobe configurations. Under each beam configuration, several symbols are captured. In other words, the beam switching rate is slower than the symbol rate. We evaluate the impact of antenna subset size, number of sidelobe configurations, and number of symbols per beam in Sec. V-C.
@@
We have implemented Panoptic using both benchtop equipment as well as high-speed RFSoCs. In the former, we use an arbitrary waveform generator and real-time scope as baseband units on the TX and RX sides, respectively. In this case, the symbol rate is MHz and our frame structure contains 20 preamble symbols followed by 1500 data symbols. The symbol rate is set to 25 MHz for one single tone. This is much higher than what standards today support. For mmWave bands, 5G NR commonly uses subcarrier spacings of 120 kHz and 240 kHz. We also implemented Panoptic programmable RFSoC FPGA as baseband units on both ends. The ADC sample rate is GHz, and the bandwidth is GHz with 192 active subcarriers. Our frame structure contains a standard IEEE 802.11 preamble and 64 data symbols.
On the TX side, the mmWave array is connected to the baseband unit, sending modulated data streams with different modulation and coding rates. On the RX side, the downconverted IQ samples are captured and processed by another baseband unit. Differential IQ ports are connected for both TX and RX to remove the distortion and noise. The TX and RX clocks are not synchronized (the nodes are not wire-connected). Panoptic’s algorithm is written in a single Python code controlling the Sivers array configuration and data capture. For all experiments, the center frequency is GHz. If not specified, we place the TX and RX 3 m apart (we evaluate the impact of TX-RX distance in Sec. V-D). We also use preambles as in a conventional communication system for packet detection. The preambles are used for coarse carrier frequency offset removal, adopting conventional signal processing schemes from the literature [43]. The data symbols are used for joint data recovery and NLOS sensing.
@@ V-A Compressive Sidelobe Radiation Patterns
Results. We measure the radiation pattern under different mainlobe directions and we randomly turn off 4 out of 16 antenna elements. The measured radiation patterns are shown in Fig. 5. We show examples of two different mainlobe configurations and two distinct random sidelobe patterns under each mainlobe. In addition, we have simulated the theoretical radiation pattern using the knowledge of antenna geometry, inter-element spacing, and randomly chosen antenna subsets in measurements. We note that the internal non-idealities between antenna elements (e.g., due to variations in the lengths of the transmission lines in the circuit board connecting the RFIC to the antennas), require a one-time calibration for the array. We have performed antenna calibration based on the beambook provided by the manufacturer, i.e., by implementing a technique known as EiCal in the literature [44]. Fig. 5 illustrates good agreement between the simulated and measured radiation patterns. Note that we collected tens of different radiation patterns overall and validated the random sidelobe generation in all cases. Fig. 5(c) and Fig. 5(d) depict a few examples of measured sidelobe-varying beam patterns while the mainlobe remains unchanged toward 0 and 45, respectively.
@@ V-C2 Number of data symbols per beam ()
Second, Fig.7(b) shows the impact of the number of data symbols captured per compressive beam at the receiver. We note that the symbol rate is always higher than the antenna switching rate in Panoptic, and hence, several symbols can be captured under a given beam configuration.Here we set the number of beams to 189. Recall from Eq. (8) that several measurements with a fixed beam help with averaging out noise and better mapping the sidelobe perturbations to the NLOS path signature. This is also reflected in our experimental results in Fig.7(b). We observe that the mean sensing error decreases with more observations per beam. With 25 symbols per beam, the mean NLOS angle error is 2.57, very close to the final value. The average reflector estimation error decreases below 2 after the symbol number exceeds 700. Also, the standard deviation decreases with more observations, which indicates a robust estimation. In Panoptic, the symbol rate does not play any major role in the sensing accuracy. Depending on the symbol rate for the communication needs, we can adapt the antenna switching rate accordingly to allow for capturing the desired number of symbols in one beam configuration.
@@ V-C3 Antenna subset Size ()
Finally, we evaluate the impact of antenna subset size, i.e., the number of randomly selected OFF antennas. To this end, we repeat the experiments by turning off 1,2,3, and 4 antennas (out of 16) while placing a reflector at various angular locations within the range of -30 to 30 degrees (relative to TX). Recall that when all antennas are ON, no sidelobe perturbation happens, and hence NLOS sensing is not possible. We note that the total number of possible sidelobe configurations is only = =16 and = =120, for the case with 1 and 2 antennas OFF, respectively. For a larger number of OFF antennas, the random beam space size is much larger than the beambook memory in the Sivers arrays, and hence we limit the number of tested beam configurations to .
@@
On the other side, when turning off more antennas, the mainlobe directivity gain is reduced, which is also captured in Fig.8(b) (all antennas ON in the leftmost). This may hinder the data detection performance with Panoptic. In this particular experiment, the BER remained at 0 (out of transmitted bits) in all cases as the gain of the mainlobe was sufficiently high to close a 3-meter link between the TX and RX even when using 12 antennas only. We evaluate the impact of TX-RX distance in Sec. V-D. We underline that the communication penalty becomes negligible with a larger antenna array because fewer antennas have to be turned OFF. For example, the randomness space size for the 16-element array with 4 antennas OFF is , and the size for a 64-element array with 2 antennas OFF is . However, turning off 2 out of 64 antennas decreases the mainlobe directivity (and consequently the SNR of the data link) roughly by dB. According to the MCS -EVM table in IEEE 802.11ad MCS [25], such an SNR reduction will not trigger an MCS downgrade.
Next, we look at the minimum antenna subset size () required to achieve an average angle error of less than 2. Our simulation considers a Tx-Rx distance of 2 m using 189 random beams. 200 randomly located reflector configurations are considered, such that they are all within +/- 45 degrees of the TX/RX nodes and the reflectors create a specular NLOS path between the TX and RX. We repeat this simulation for different array sizes from 12 to 64. The results are shown in Fig. 9. We observe that the number of antennas OFF required for maintaining the same sensing accuracy slightly decreases with a larger array size. For example, a 12-element array requires 4 antennas to be turned off, but a 64-element array only needs 2. We also note that the antenna array with 8 or 10 antennas fails to achieve the sensing resolution even when turning off half of its elements (maximizing the randomness space). The best achievable angle estimation errors are 4.12° and 2.69°, respectively. In practice, the antenna subset size can be adapted in Panoptic to balance between angular localization accuracy and the desired BER.
@@ V-D4 Mobile Reflector Trajectory
So far, we have configured a reflector moving perpendicular to the TX-RX LOS path, similar to Trace 1 in Fig. 13(a). Here, we argue that Panoptic is robust to other arbitrary trajectories that a reflector might take as long as the object’s reflection footprint can be observed at the receiver. Hence, we place the TX and RX at a 3 m distance in a conference room and transmit 4QAM data symbols at a rate of 25 MHz. We configure three different trajectories as shown in Fig. 13(a). The estimation results are depicted in Fig. 13(b), 13(c) and 13(d), respectively. We observe that the ground truth angles match well with angles estimated by Panoptic. The error is higher around 0 for all three traces as the reflector is partially blocking the LOS path. Further, the error increases as the reflector approaches the FoV of TX or RX antenna arrays. Overall, we observe that the NLOS sensing accuracy is oblivious to the distance of the reflector from the TX or RX.
@@
Finally, it is worth mentioning that while we adopt a relatively low antenna switching rate for proof of concept experiments, it is possible to increase the antenna switching rate in practice for faster environmental tracking. With a simple calculation, we find that the antenna switching rate needed to angularly track mobile targets is well below the capability of today’s off-the-shelf arrays, allowing for real-time estimation and angular tracking in practical mobile environments. Specifically, given a symbol rate of and the number of data symbols per beam , the beam switching rate can be found as . For instance, with a symbol rate of GHz and minimum required (for reliable NLOS signature extraction as discussed in Sec. V-D), we reach the beam switching rate of MHz, which is an order of magnitude smaller than the Sivers array switching rate of MHz. Indeed, previous work showed an antenna switching rate of about MHz on Sivers arrays with FPGA [16]. Note that supporting higher symbol rates is trivial, as one can consider increasing , which helps with better averaging and NLOS signature extraction from weak sidelobe perturbations. Further, as shown by our experimental results in Sec. V-D, testing with different random sidelobe configurations is sufficient for accurate NLOS sensing. Hence, the sensing frequency is calculated by KHz, following the previous example. Such a high sensing rate is sufficient to angularly track even high-speed targets in dynamic environments, e.g., cars. Specifically, at this rate, even a fast-moving object such as a car traveling at 100 km/h (27.8 m/s) moves only about 2.1 mm during the sensing intervals, which is negligible and would not yield any meaningful AOA variations.
V-D5 Panoptic Implementation on TX-Side vs RX-side
Here, we attempt to answer a simple question: What if Panoptic is adopted at the transmitting beams instead of receiving beams? To answer this question, we conduct simulations in which we consider different array sizes on one end (TX or RX) while the other end has a simple omnidirectional beam. We set 189 random compressive sidelobe beams with 4 antennas turned off, regardless of the array size. The distance between TX and RX is set at 2 meters. 50 randomly located reflector configurations are considered, such that they are all within +/- 45 degrees of the TX/RX nodes and the reflectors create a specular NLOS path between the TX and RX. The NLOS angle estimation result is shown in Fig. 14. We observe that Panoptic’s sensing performance is not a function of the TX or adoption, but instead, it depends on the number of random beam configurations that are tested. The slight improvement with larger arrays comes from a greater number of distinct sidelobe patterns. As array size increases, the space of possible OFF positions expands, reducing sidelobe correlation and improving compressive sensing performance.
V-D6 Impact of the TX Beamwidth on Communication and Sensing Performance
We also explore how the TX beamwidth affects both communication and sensing performance, assuming a fixed TX beam and the Rx adopting compressive sidelobe beams. In our simulation, we vary the TX beamwidth from 1 to 180, and use a 16-element Rx antenna array with 4 elements turned OFF. We generate 189 random compressive sidelobe beams, with the TX and RX spaced 2 m apart. We consider 200 random reflector locations, all within of the nodes. Each reflector creates a specular NLOS path between the TX and RX. The result is shown in Figs. 15(a) and 15(b). As the Tx beam becomes wider, the communication power decreases because the energy is spread over a broader area. On the other hand, the NLOS sensing accuracy improves since more of the surrounding space is illuminated, and reflections from the environment are more likely to be captured. In practice, an adaptive algorithm can be used to adjust the TX beamwidth depending on different application scenarios that vary in the required link capacity, SNR range, environment complexity, etc., balancing between communication and sensing performance.
@@ VI Discussion
Super-Resolution Sensing with Panoptic. While this paper builds a novel foundation for joint communication and sensing using antenna subset switching, there are several schemes that can be adopted in future work to improve the sensing resolution. First, combining phase and amplitude readings, one can extend Panoptic to more advanced sensing tasks like environmental imaging, human activity recognition, etc. We leave this exploration for future work. Doppler extraction, which offers information about the velocity of objects in the environment, is left for future work. Extracting Doppler shifts requires coherent phase measurements across successive transmissions, which in turn demands tighter synchronization between nodes and stable local oscillators—challenges that are more pronounced in mmWave systems. With Doppler information, our system could capture dynamic environmental changes, enabling several new applications.
Second, Panoptic chooses the antenna subset randomly without any feedback from the RX. In principle, not every sidelobe perturbation pattern provides the same amount of information for a particular target extraction. In the future, we will explore adaptive compressive sidelobe selection strategies. Last but not least, more advanced data processing methods, such as data-driven and ML-assisted techniques, could potentially improve sensing accuracy. Machine learning-assisted methods hold strong potential for improving sensing accuracy due to their ability to capture nonlinear features and robustness to noise. In the current approach, Panoptic builds a simulated library of NLOS signatures and uses correlation to estimate angles by matching received signatures to this library. This process could be enhanced using sequence-based machine learning models. Specifically, the sequence of measurements from different compressive sidelobe beams can be treated as input, with the goal of predicting the angle of arrival. Well-established models like RNNs or Transformers are well-suited for this task. A practical strategy would involve pretraining these models on simulated data to capture the structure of Panoptic ’s algorithm, followed by fine-tuning on a small set of real-world measurements to adapt to hardware imperfections and environmental nuances. We leave this for future work.
Scaling to Large Arrays . The mmWave nodes we use in evaluating Panoptic have 16 antenna elements on the azimuth plane. Indeed, mmWave AP/BS (and even mmWave handheld devices) are often equipped with larger antenna arrays. Larger antenna arrays expand the randomness space for sidelobe perturbations and result in better angular localization resolutions. In this case, turning off a small number of antennas can provide sufficient random space for sensing; hence, the penalty of compressive sidelobe transmission on the communication SNR and BER becomes negligible. As discussed in Sec. III-C2, the number of compressive sidelobe beam measurements only grows with , where is the number of antennas in array; hence, Panoptic is scalable to large arrays.